# RF-GML: Reference-Free Generative Machine Listener

### v1

@@ I Introduction
Most RF models are specifically designed for speech, typically at a 16 kHz sample rate, and focus on estimating quality degraded by noise or enhanced by deep noise suppression methods. Specifically for coded speech, [11] introduced an RF model to estimate PESQ (an FR metric) scores from various wideband codecs at a 16 kHz sample rate. WaweNet [6] and NISQA [7] can also predict the quality of coded speech, with NISQA predicting transmitted speech quality at a 48 kHz sample rate. Recently, SESQA [8], an RF metric for speech at a 48 kHz sample rate, was proposed, improving accuracy over NISQA and other RF metrics by training on multiple objectives, such as Mean Opinion Score (MOS), ranking, and FR metrics.
@@ III-A Training set
We used our internal MUSHRA listening tests at a 48 kHz sample rate, all conducted with headphones. Each test featured an unencoded hidden reference, 3.5 kHz and 7 kHz low-pass filtered versions of the reference, and one or more coded signals. The stereo codecs in the training set included AAC [22], HE-AAC v1 and v2 [23], and Dolby AC-4 [24], covering various bitrates and all content types.
@@ IV-A Model architecture
Like GML, RF-GML uses an inception-based convolutional model for stereo [14] as the backbone, with architectural details provided in Fig. 3 and Table 1 of [14], and design motivation discussed in [13]. In short, the model combines four modified Inception blocks (In-A, In-A, In-B, In-C) adapted for audio, interspersed with three Squeeze-and-Excitation (SE) blocks, followed by three fully connected layers. The first two Inception blocks share the same kernel shapes, so both are named In-A. Unlike GML, RF-GML processes only Gammatone spectrograms for the degraded left ($L$), right ($R$), mid ($M=(L+R)/2$), and side ($S=(L-R)/2$) signals. The output stage, similar to GML, is adapted to provide two parameters describing the distribution of MUSHRA scores, as outlined in Section II. As with GML, individual listener scores are used for training rather than mean subjective scores. Our proposed RF-GML has 15.25M parameters, 0.015% fewer than GML. Since GML likely learned useful weights for predicting coded audio quality, we propose transferring the weights of degraded input channels from the first In-A block in GML and retraining with listening tests. The benefits of the transfer learning approach are demonstrated in Section IV-C.
@@ IV-B Training configuration
The training dataset was normalized and randomly split into 80% for training and 20% for validation. To fully utilize the listening scores, we applied 5-fold cross-validation. The model was implemented in PyTorch and trained for 10 epochs per fold (50 epochs in total) on an Nvidia V100 GPU using the Adam optimizer. We used the optimal kernel sizes from [13],[14], a learning rate of $10^{-4}$, and a batch size of 8, following the initialization scheme mentioned in Section II. We also used CutMix as a data augmentation strategy, which randomly cuts and pastes spectrogram regions to create new spectrograms with associated quality scores $\tilde{y}$:
@@ IV-C RF-GML benchmarking and evaluation
Figure 1 shows that GML has the lowest NLL loss and is expected to have the highest prediction accuracy. Table 1 further confirms its superior accuracy over ViSQOL [15]. While RF-GML (def) and RF-GML (deg) have similar training losses, Table 1 indicates that RF-GML (deg) performs best across all test sets and metrics, often achieving top-1 scores and reliably rating unencoded audio closer to 100. We speculate this is due to knowledge transfer from GML, which was trained on many reference-reference signal pairs. Although RF-GML (deg) may have lower correlation scores at times, it still consistently ranks higher than RF-GML (def), which performs better in correlation metrics but overall is second best. RF-GML (degF) shows decreased prediction accuracy when initialized like RF-GML (def) but frozen, underscoring the need for both initialization and learning from the first inception (In-A) block. Simply initializing all inception block weights from GML further degrades unencoded reference ratings. Due to slightly poorer NLL losses, both RF-GML (degF) and RF-GML (all) are excluded from further analysis. Overall, considering all five listening test sets together (evaluating 953 signals), RF-GML (deg) achieves $R_{p}$, $R_{s}$, and MU of 0.83, 0.83, and 89.56, respectively, compared to the second-best RF-GML (def), which achieves 0.83, 0.81, and 82.14, respectively. Table 1 shows that, for the mono and stereo tests, these two RF-GML models achieve a correlation with subjective ratings comparable to ViSQOL.
@@
Next, we encoded all 24 stereo excerpts from the test set at various bitrates using the HE-AAC codec family commonly applied in practice, and demonstrated (Figure 2) that the mean RF-GML quality score scales well with bitrate. However, we notice that the average quality predicted with RF-GML (deg) at 96 kb/s HE-AAC v1 is slightly higher than that of AAC at 192 kb/s. A possible reason could be that HE-AAC v1 at 96 kb/s has a higher overall bandwidth due to the use of spectral band replication [23]. Since we would expect the quality of both 192 kb/s AAC and 96 kb/s HE-AAC to fall in the excellent quality range, this would have been a difficult task even for human evaluation in a reference-free quality assessment. On the other hand, RF-GML (def) predictions saturate around 80, with less quality separation across bitrates and a slight quality increase when dropping from 32 to 20 kb/s. This observation, along with strong performance in predicting subjective quality (as shown in Table I), makes RF-GML (deg) more suitable for quality monitoring in adaptive streaming clients.
@@
Finally, we evaluated the model’s performance in rating unencoded content by running RF-GML and SESQA on 511 excerpts, including synthetic test tones and signals, and plotted scatter plots of predicted quality scores versus bandwidth (Figure 3). This content, curated internally over the years for testing the engineering implementation of codecs, was not fully known to the authors before the experiments. Figure 3 shows almost no correlation (-0.08) between bandwidth and quality scores for SESQA, while RF-GML (deg) and RF-GML (def) show slight correlations of 0.44 and 0.51, respectively. This is expected, as RF-GML is trained to evaluate coding artifacts, with audio bandwidth being a typical audio codec tuning parameter. Note that a better correlation with RF-GML (def) may not necessarily indicate better performance because unencoded speech signals may also have lower bandwidth. The scatter plot again shows that RF-GML (deg) rates unencoded audio closer to 100 than RF-GML (def). The five lowest-rated signals for RF-GML are the 3.5 kHz and 4.5 kHz low-pass filtered versions, while SESQA scored full-bandwidth unencoded music signals the lowest. This is expected for RF-GML, which was trained on MUSHRA tests with low-pass filtered anchors, and for SESQA, which was not trained with music signals. Both SESQA and RF-GML (deg) rated the five lowest bandwidth signals (1 kHz sine tones and an $\approx$500 Hz tone from a gong instrument) similarly, assigning high-quality scores. These results demonstrate that RF-GML (deg) consistently rates unencoded audio closer to 100 and could potentially be used to identify unencoded audio.
@@ V Conclusions
This paper presents RF-GML, a novel reference-free audio quality metric capable of evaluating coded mono, stereo, and binaural audio at 48 kHz. Extensive benchmarking has demonstrated RF-GML’s superior performance, particularly in accurately rating unencoded audio and distinguishing between different levels of coding artifacts. The transfer learning approach, which leverages knowledge from a pre-trained full-reference GML, has proven effective in enhancing RF-GML’s performance. By initializing the weights of degraded input channels with corresponding weights from the full-reference model, RF-GML gains a better understanding of reference signals, leading to more accurate quality predictions. The proposed metric offers several advantages: it is versatile and capable of evaluating a wide range of audio content and coding scenarios. Additionally, RF-GML’s ability to accurately rate unencoded audio is valuable for identifying high-quality content in large-scale music distribution and archiving. While RF-GML can predict confidence intervals, this aspect has not been evaluated in the paper and will be explored in future work following the construction of reference-free subjective tests.

### v2

@@ I Introduction
While most RF models are designed for speech, typically at a 16 kHz sample rate, and focus on estimating quality degraded by noise or enhanced by deep noise suppression methods, some models address a broader range of impairments. Specifically for evaluating coded speech, [11] introduced an RF model to estimate PESQ (an FR metric) scores from various wideband codecs at a 16 kHz sample rate. WaweNet [6] and NISQA [7] can also predict the quality of coded speech, with NISQA predicting transmitted speech quality at a 48 kHz sample rate. Recently, SESQA [8], an RF metric for speech at a 48 kHz sample rate, was proposed, improving accuracy over NISQA and other RF metrics by training on multiple objectives, such as Mean Opinion Score (MOS), ranking, and FR metrics.
@@ III-A Training set
We conducted our internal MUSHRA listening tests at a 48 kHz sample rate using headphones. Each test featured an unencoded hidden reference, 3.5 kHz and 7 kHz low-pass filtered versions of the reference, and one or more coded signals. The stereo codecs in the training set included AAC [22], HE-AAC v1 and v2 [23], and Dolby AC-4 [24], covering various bitrates and all content types.
@@ IV-A Model architecture
Like GML, RF-GML uses an inception-based convolutional model for stereo [14] as the backbone, with architectural details provided in Fig. 3 and Table 1 of [14], and design motivation discussed in [13]. In short, the model combines four modified Inception blocks (In-A, In-A, In-B, In-C) adapted for audio, interspersed with three Squeeze-and-Excitation (SE) blocks, followed by three fully connected layers. The first two Inception blocks share the same kernel shapes, so both are named In-A. Unlike GML, RF-GML processes only Gammatone spectrograms for the degraded left ($L$), right ($R$), mid ($M=(L+R)/2$), and side ($S=(L-R)/2$) signals. The output stage, similar to GML, is adapted to provide two parameters describing the distribution of MUSHRA scores, as outlined in Section II. As with GML, individual listener scores are used for training rather than mean subjective scores. Our proposed RF-GML has 15.25M parameters. Since GML likely learned useful weights for predicting coded audio quality, we propose transferring the weights of degraded input channels from the first In-A block in GML and retraining with listening tests. The benefits of the transfer learning approach are demonstrated in Section IV-C.
@@ IV-B Training configuration
The training dataset was normalized and randomly split into 80% for training and 20% for validation. To fully utilize the listening scores, we applied 5-fold cross-validation. The model was implemented in PyTorch and trained for 10 epochs per fold (50 epochs in total) on an Nvidia V100 GPU using the Adam optimizer. We used the optimal kernel sizes from [13],[14], a learning rate of $10^{-4}$, and a batch size of 8, following the initialization scheme described in Sections II and IV-C. We also used CutMix as a data augmentation strategy, which randomly cuts and pastes spectrogram regions to create new spectrograms with associated quality scores $\tilde{y}$:
@@ IV-C RF-GML benchmarking and evaluation
Figure 1 shows that GML has the lowest NLL loss and is expected to have the highest prediction accuracy. Table 1 further confirms its superior accuracy over ViSQOL [15]. While RF-GML (def) and RF-GML (deg) have similar training losses, Table 1 indicates that RF-GML (deg) performs best across all test sets and metrics, often achieving top-1 scores and reliably rating unencoded audio closer to 100. We speculate this is due to knowledge transfer from GML, which was trained on many reference-reference signal pairs. Although RF-GML (deg) may have lower correlation scores at times compared to RF-GML (def), considering all the test sets and metrics, it outperforms RF-GML (def). RF-GML (degF) shows decreased prediction accuracy when initialized like RF-GML (def) but frozen, underscoring the need for both initialization and learning from the first inception (In-A) block. Simply initializing all inception block weights from GML further degrades unencoded reference ratings. Due to slightly poorer NLL losses, both RF-GML (degF) and RF-GML (all) are excluded from further analysis. On average, considering all five listening test sets together (evaluating 953 signals), RF-GML (deg) achieves $R_{p}$, $R_{s}$, and MU of 0.83, 0.83, and 89.56, respectively, compared to the second-best RF-GML (def), which achieves 0.83, 0.81, and 82.14, respectively. Table 1 shows that, for the mono and stereo tests, these two RF-GML models achieve a correlation with subjective ratings comparable to ViSQOL.
@@
Next, we encoded all 24 stereo excerpts from the test set at various bitrates using the HE-AAC codec family commonly applied in practice, and demonstrated (Figure 2) that the mean RF-GML quality score scales well with bitrate. However, we notice that the average quality predicted with RF-GML (deg) at 96 kb/s HE-AAC v1 is slightly higher than that of AAC at 192 kb/s. A possible reason could be that HE-AAC v1 at 96 kb/s has a higher overall bandwidth due to the use of spectral band replication [23]. Since we would expect the quality of both 192 kb/s AAC and 96 kb/s HE-AAC to fall in the excellent quality range, this would have been a difficult task even for human evaluation in a reference-free quality assessment. On the other hand, RF-GML (def) predictions saturate around 80 on this data, with less quality separation across bitrates and a slight quality increase when dropping from 32 to 20 kb/s. This observation, along with strong performance in predicting subjective quality (as shown in Table I), makes RF-GML (deg) more suitable for quality monitoring in adaptive streaming clients.
@@
Finally, we evaluated the model’s performance in rating unencoded content by running RF-GML and SESQA on 511 excerpts, including synthetic test tones and signals, and plotted predicted quality scores versus signal bandwidth in scatter plots (Figure 3). This content, curated internally over the years for testing the engineering implementation of codecs, was not fully known to the authors before the experiments. Figure 3 shows almost no correlation (-0.08) between bandwidth and quality scores for SESQA, while RF-GML (deg) and RF-GML (def) show slight correlations of 0.44 and 0.51, respectively. It is evident that SESQA is unsuitable for rating unencoded content. A higher correlation with bandwidth is expected with the RF-GML variants, as they are trained to evaluate coding artifacts, with audio bandwidth being a typical codec tuning parameter across bitrates (or quality levels). However, note that a better correlation with RF-GML (def) may not necessarily indicate better performance, as unencoded speech signals may also have lower bandwidth (e.g., 16 kHz), though ideally, the predicted quality should be close to 100. The scatter plot further shows that RF-GML (deg) rates unencoded audio closer to 100 than RF-GML (def). The five lowest-rated signals for RF-GML are the 3.5 kHz and 4.5 kHz low-pass filtered versions, while SESQA scored full-bandwidth unencoded music signals the lowest. This is expected for RF-GML, which was trained on MUSHRA tests with low-pass filtered anchors, and for SESQA, which was not trained with music signals. Both SESQA and RF-GML (deg) rated the five lowest bandwidth signals (1 kHz sine tones and an $\approx$500 Hz tone from a gong instrument) similarly, assigning high-quality scores. These results demonstrate that RF-GML (deg) consistently rates unencoded audio closer to 100.
@@ V Conclusions
This paper presents RF-GML, a novel reference-free audio quality metric capable of evaluating coded mono, stereo, and binaural audio at 48 kHz. Extensive benchmarking demonstrates RF-GML’s superior performance. The transfer learning approach, which leverages knowledge from a pre-trained full-reference GML, has proven effective in enhancing RF-GML’s accuracy. By initializing the weights of degraded input channels with corresponding weights from the full-reference model, RF-GML gains a better understanding of reference signals, leading to more precise quality predictions. The proposed metric offers several advantages: it is versatile and capable of evaluating a wide range of audio content and coding scenarios. While RF-GML can predict confidence intervals, this aspect has not been evaluated in this paper and will be explored in future work, following the construction of reference-free subjective tests.